# Grounding Synthetic Data Evaluations of Language Models in Unsupervised Document Corpora

### v1

@@ 5.1.3 Impact of Question Length
Its worth noting that the same effect does not show up in answer length. More detailed (longer) MCQ answer options does not produce the same increase in benchmark evaluation accuracy as longer questions. Figure 6 demonstrates this via scatterplot showing synthetic benchmark accuracy over-performance (similar to Figure 5). Only the gpt-4.1-mini at the very right of the plot significantly is biased above the $y=0$ expected trend line.
@@
The following question answer pair was generated by gpt-4.1-mini (37 tokens) using a context from Squadv2, highlighting the difference in model verbosity:
@@
6 "D": "Overuse of antibiotics in humans alone causes resistance, while use in livestock has no significant impact on bacterial evolution or public health."
7},
8"answer": "A"
Contrasted with this question answer pair generated by Llama-4-Maverick-17B-128E-Instruct using the same context and topic, producing a significantly shorter question (22 tokens) and answers:
@@
6 "D": "By inducing genetic mutations in human pathogens that confer resistance."
7},
8"answer": "B"
5.2 Synthetic Benchmark Question Diversity
During synthetic question generation, the explicit topic extraction step provides a backstop against degenerate questions which might focus only on the most obvious elements of a context chunk. To compare the variety and coverage of the generated questions we compute the average cosine similarity (using sentence-transformers/all-MiniLM-L6-v2 as the embedding model) between all question pairs (excluding self pair) over the generated synthetic dataset. Figure 7 shows the question diversity per generating model along with the original question diversity.
Figure 7: Synthetic question diversity (pairwise cosine distance using sentence-transformers/all-MiniLM-L6-v2) per generating model. Violin bars show the spread across the 5 datasets which have original questions.
@@ 5.2 Synthetic Benchmark Question Diversity
It is interesting to note that some highly capable models (such as gpt-4.1) show comparatively worse performance on generated question diversity, when the same model performs well in clarity, correctness, and groundedness.
5.3 Benchmark Accuracy Correlation
Ideally, we want synthetic benchmarks to correlate well with the evaluation accuracy measured by human curated questions. This ends up being true for some generating models, but not all. Figure 8 shows the evaluation accuracy scatterplot comparing the human questions against the synthetic generated benchmark questions for the 5 bolded ensemble models in section 4.2. These models end up producing a Pearson accuracy correlation of $0.745$ ($R^{2}=0.555$) for the evaluation accuracies for synthetic vs human benchmark questions. Each plot marker is the average evaluation accuracy for a given model-dataset combination.
Figure 8: MCQ synthetic vs human evaluation accuracy exploring the Pearson correlation in benchmark performance for the ensemble of five generating models.
There is a sec_qa dataset specific oddity: The over-performance on human questions (top right) stems from the multiple choice human questions being very easy, and producing an approximately flat line over various synthetic versions generated by LMs. Interestingly enough, this over-performance disappears under the open-ended question regime for this dataset as seen in fig. 13. However, while this ensemble of models produces reasonable overall Pearson accuracy correlation between human and synthetic questions, occasionally specific LMs generate undesirable benchmarks. Figure 9 shows two example benchmark evaluation comparisons, the left plot (a) where the synthetic benchmark shows reasonable evaluation quality with Pearson accuracy correlation between human and synthetic questions generated just by Llama-4-Maverick-17B-128E-Instruct, and the right (b) where the synthetic benchmark generated by gpt-4.1-mini does not (the questions are too easy).
(a) Generating Model Llama-4-Maverick-17B-128E-Instruct
(b) Generating Model gpt-4.1-mini
Figure 9: MCQ synthetic vs human benchmark evaluation performance exploring Pearson accuracy correlation. (a) reasonable Pearson capability correlation between human and synthetic. (b) significant over-performance on synthetic questions, highlighting overly easy generated questions.
5.4 Benchmark Model Ranking Correlation
While exact Pearson correlation between human and synthetic benchmark evaluation accuracy is a nice to have feature, evaluations are used to compare models. Therefore, a more important metric is how well does the rank order of models correlate when comparing benchmark evaluation accuracy. Spearman rank correlation measures this ranking stability. Figure 10 plots the evaluation model performance across all datasets, comparing the synthetic and human accuracy Spearman ranking correlation of the evaluation model capability. This matches the use case of a user selecting between various models for a specific application relevant to the provided grounding documents.
(a) Benchmark evaluation accuracy generated by Llama-4-Maverick-17B-128E-Instruct.
(b) Benchmark evaluation accuracy generated by gpt-4.1-mini.
Figure 10: Comparison of benchmark rank ordering between synthetic and human evaluation accuracy. (a) Questions demonstrate acceptable alignment and Spearman rank correlation $0.954$. (b) Question generation failure, showing gpt-4.1-mini over-estimating LM capability via overly easy generated questions (Spearman rank correlation $0.607$).
Figure 10 (a) shows the Spearman rank order equivalent of Figure 9 (a) which scatterplots the evaluation accuracy and Pearson accuracy correlation. Overall Pearson evaluation accuracy correlation translates into Spearman ranking correlation and relative model order stability. Similarly, the overly easy synthetic benchmark in Figure 9 (b) produces a mis-matched rank ordering (Spearman correlation $0.607$) in Figure 10 (b). While ensembling various question generating LMs together will mitigate this effect, it’s worth validating the set of generating models against the datasets in this paper to confirm the ensemble is not decreasing the overall synthetic benchmark quality. The tendency of gpt-4.1 and gpt-4.1-mini to produce overly long detailed questions that LMs find easier to answer might be the cause of the observed benchmark over-performance.
5.5 Multiple-Choice Questions vs Open-Ended
@@
During this study both multiple choice and open-ended questions were evaluated. None of our major findings change whether using multiple choice questions for simplified evaluation benchmark grading, or open-ended questions to produce harder evaluation benchmarks where knowledge synthesis is required. The models good at creating synthetic multiple choice questions are also good at creating open ended questions. The models which create overly easy questions by including too much detail (like phi-4) do so for both open ended and MCQ, with the long question problem becoming even more pronounced for open ended questions as seen in Figure 11.
(a) Generated question token count by generating model for open-ended questions.
@@
(b) Impact of question token count on the benchmark evaluation accuracy over-performance.
@@
Figure 11: Longer synthetic open-ended questions have an impact on evaluation accuracy over-performance. (a) Open-ended questions demonstrate an exacerbated question length problem compared to MCQ. (b) Longer synthetic open-ended questions correlate with over-performance on the evaluation benchmark.
@@
Table 2 showcases the Spearman rank correlation between evaluation the benchmarks written by humans and the synthetic data versions. The table includes data for all three of the human question reformatting models used to demonstrate the minimal impact that reformatting has on the overall high level results. Overall, the best outcome is to ensemble together a few questions generated by various reasonably capable LMs to avoid the known slight self-bais/self-preference issue. While we used Llama-4-Maverick as the text feature extractor for measures like groundedness and clarity, that model was unexpectedly dominant in these final results.
@@
Generating Model OE Rank MC Rank OE Rank MC Rank OE Rank MC Rank
Ensemble 0.9648 0.9679 0.9516 0.9571 0.9209 0.9821
Llama-3.2-3B-Instruct 0.7143 0.6143 0.7143 0.6929 0.6791 0.6536
Llama-3.3-70B-Instruct 0.8813 0.9571 0.8681 0.9643 0.8242 0.9714
Llama-3.1-8B-Instruct 0.7275 0.7964 0.8110 0.7321 0.7275 0.7714
Llama-4-Maverick 0.9648 0.9679 0.9516 0.9536 0.9209 0.9821
gemma-3-12b-it 0.8198 0.8929 0.8418 0.8964 0.7890 0.9000
gemma-3-27b-it 0.9516 0.9571 0.9472 0.9857 0.9121 0.9714
gemma-3-4b-it 0.9209 0.6393 0.8989 0.6571 0.8637 0.6857
gpt-4.1 0.8549 0.9179 0.8725 0.9250 0.8110 0.9357
gpt-4.1-mini 0.8418 0.5643 0.8549 0.6286 0.7934 0.6071
o4-mini 0.7934 0.9250 0.8374 0.9036 0.7714 0.9429
phi-4 0.1780 0.6964 0.3010 0.6250 0.1736 0.7607
Qwen2.5-7B-Instruct 0.8901 0.7321 0.8945 0.7143 0.8505 0.7786
Table 2: Spearman rank correlation between synthetic and human benchmarks for open-ended (OE) and multiple choice question (MCQ) evaluations across generating models. Ensemble = bolded models.
Table 3 contains the benchmark evaluation Pearson accuracy correlation between the reformatted human questions and the synthetic data generated questions. Similarly to the Spearman ranking, Llama-4-Maverick performed very well in these tests. However, unlike the ranking data, using a ensemble of models produces significantly worse Pearson accuracy correlation numbers, largely due to the long and overly easy questions written by gpt-4.1 and o4-mini.
@@
Generating Model OE Corr MC Corr OE Corr MCQ Corr OE Corr MCQ Corr
Ensemble 0.6437 0.4752 0.7248 0.5630 0.6569 0.4798
@@
Table 3: Pearson correlation of evaluation accuracy between synthetic and human benchmarks for open-ended (OE) and multiple-choice (MC) evaluations across generating models. Ensemble = bolded models.
5.6 Case Study on Two Academic Papers
To demonstrate our approach in a more realistic setting, two academic papers were selected: “Current Solutions and Future Trends for Robotic Prosthetic Hands” (labeled annurev-control-071020-104336[55] and “Recent Advances in Large Language Model Benchmarks against Data Contamination: From Static to Dynamic Evaluation” (labeled arXiv_2502_17521v1[56]. The documents were chunked into 27 and 28 contexts respectively. “Current Solutions” averaged $7.7$ topics per context for a total of 210 question topics. Topic extraction required $51k$ input and $19k$ output tokens. “Recent Advances” averaged $6.9$ topics per context for a total of 194 question topics. Topic extraction required $34k$ input and $15k$ output tokens. Figure 12 showcases the evaluation performance of LMs on our two datasets, where the MCQ results show most models performing well before a sharp dropoff with 4B and fewer parameter models, and the Open-Ended results indicating a much smoother decline in performance. Multiple-choice questions can be easier to answer due to the presence of the answer options, even if the evaluation framework shuffles the option order. This effect is likely what is causing the MCQ evaluation accuracy to remain higher for longer.
@@ 5.6 Case Study on Two Academic Papers
Figure 12: MCQ (a) Compared to Open-Ended (b) Synthetic evaluation benchmark results for two papers drawn from current research work ([55, 56].
The following example questions were drawn from the MCQ synthetic benchmark from arXiv_2502_17521v1 [56].
@@
6 "D": "Temporal cutoff generates data points using LLMs, while rule-based generation uses manual curation."
7},
8"answer": "A",},
9
10{"question": "What is a potential future research direction for improving the design and standardization of dynamic benchmarking methods for Large Language Models (LLMs), given the current limitations and proposed evaluation criteria?",
11"choices": {
@@
15 "D": "Focusing solely on post-hoc contamination detection methods."
16},
17"answer": "A",}
@@
2 "answer": "By leading to misleading conclusions that favor contaminated models.",},
3
4{"question": "What are the implications of using LLMs to rewrite benchmark samples while preserving their original difficulty levels, as seen in methodologies like ITD, on the overall diversity and complexity of the resulting benchmark datasets?",
5 "answer": "The resulting datasets may maintain complexity but risk lacking diversity if not combined with other rewriting strategies.",}
5.7 Limitations
This work relies on a simple and possibly non-optimal document chunking approach. Improvements and extensions which cluster related topics or provide more fine-grained control over the information present in each chunk could improve the granularity of the generated benchmark. Including information found in tables, charts, and images would provide a richer source of information that benchmark creation could take advantage of. Additionally, improving the nuance and difficulty of the questions would provide significantly improved LM capability measurement. This likely requires careful management of the context information provided to the generating LM which is writing questions. Question generation prompting could likely be improved to encourage both more broad information synthesis and detailed, nuanced questions. Finally, the datasets used in this study to validate correlation with human written questions are all NLP extractive question answering type datasets. This limits the quality of the synthetic benchmarks as the source context material is fairly simplistic, especially compared to the two academic papers used in Section 5.6. Only the sec_qa dataset which is grounded in an open source computer security textbook, comes close to the topic diversity and complexity of the example academic papers. All other validation datasets rely upon more general trivia type knowledge.
6 Conclusion
We introduce a methodology for creating synthetic benchmarks to evaluate language model (LM) capabilities grounded in high-quality document corpora selected by practitioners. This ensures the generated benchmarks are directly relevant to their specific domains and interests. This approach extends general-purpose LM benchmarking into specialized domains, allowing for the identification of strengths and weaknesses that may not be apparent in broader evaluations.
Our method supports generating both multiple-choice questions (optimizing for ease of grading) and open-ended questions (optimizing for diagnostic relevance), allowing users to choose based on their evaluation goals. Validation against five human curated NLP dataset achieves a Spearman rank correlation of $0.96$ and a benchmark Pearson accuracy correlation of $0.79$ between synthetic and human curated questions. Additionally, we identified a potential synthetic data generation failure mode: excessively long and detailed questions can artificially inflate model performance, likely by overspecifying information related to the answer. This methodology enables the creation of targeted, relevant benchmark evaluation of domain-specific Language Model capability, facilitating deeper exploration into the capabilities and limitations of Language Models deployed for complex knowledge work.
Appendix A Multiple-Choice Questions vs Open-Ended
Figure 13 plots the correlation in evaluation accuracy between the human reformatted questions and the synthetic generated benchmark.
Figure 13: Multiple Choice vs Open Ended synthetic vs human benchmark evaluation performance exploring Pearson accuracy correlation. (a) Multiple choice Pearson accuracy correlation between human and synthetic. (b) Open-ended Pearson accuracy correlation. Both results demonstrate reasonable Pearson correlations between the benchmark evaluation accuracy.

### v2

@@ Appendix B Impact of question and answer length
While longer more detailed questions (both MC and OE) produce a model over-performance (see Figure 5), the same trend does not appear in answer length. Figure 7 demonstrates this via scatterplot showing synthetic benchmark accuracy over-performance. Only the gpt-4.1-mini at the very right of the plot is significantly biased above the $y=0$ expected trend line. The over-performance trend shown for questions in Figure 5(b) is absent.
@@
The model gpt-4.1-mini demonstrates the largest over-performance for both question length and answer length as shown in Figure 5(b). The following question answer pair was generated by gpt-4.1-mini (37 tokens) using a context from Squadv2, highlighting the difference in model verbosity in both question and answers:
@@
7 "D": "Overuse of antibiotics in humans alone causes resistance, while use in livestock has no significant impact on bacterial evolution or public health."},
8
9"answer": "A"
Contrast with this question answer pair generated by Llama-4-Maverick-17B-128E-Instruct using the same context and topic. Maverick produces a significantly shorter question (22 tokens) and answers:
@@
7 "D": "By inducing genetic mutations in human pathogens that confer resistance."},
8
9"answer": "B"
Appendix C Synthetic benchmark question diversity
@@
Most LMs produce similar diversity scores that are closely aligned with the human questions Some highly capable models (such as gpt-4.1) show comparatively worse question diversity while performing well in clarity, correctness, and groundedness. Figure 8 shows the question diversity per generating model along with the original question diversity.
Figure 8: Synthetic question diversity (pairwise cosine distance) per generating model. Violin bars show the spread across the 6 NLP datasets.
Appendix D Multiple-choice vs open-ended questions
@@
The generating models which create overly easy questions by including too much detail (like phi-4 and gpt-4.1-mini) do so for both open ended and multiple choice questions. The long question problem becomes even more pronounced for open ended questions as shown in Figure 9.
(a) MC question token count by generating model.
(b) OE question token count by generating model.
Figure 9: Most generating models are within $20\%$ of human question token counts, but some produce overly long questions. The opened-ended questions trend toward longer questions overall.
[…]
@@
(a) Impact of multiple-choice question token count on evaluation accuracy over-performance, caused by overly easy questions.
(b) Impact of open-ended question token count on evaluation accuracy over-performance, caused by overly easy questions.
@@
Figure 10: Longer synthetic open-ended questions have the same impact on evaluation accuracy over-performance as MCQ.
While similar trends in question length and the over-performance that induces, the evaluation accuracy correlations and model rank order work equally well for open ended and multiple choice questions. Figure 11(a) plots the correlation in evaluation accuracy between the human reformatted questions and the synthetic generated benchmark for multiple choice, and Figure 11(b) for open-ended. Figure 11(a) shows a sec_qa($\diamond$) dataset specific oddity: the over-performance on human questions (top right) stems from the multiple choice human questions being very easy and produces an approximately flat line over various synthetic versions generated by LMs. Interestingly enough, this over-performance disappears under the open-ended question regime for this dataset as seen in Figure 11(b).
Figure 11: Multiple Choice vs Open Ended synthetic vs human benchmark evaluation performance exploring Pearson accuracy correlation. Both results demonstrate reasonable Pearson correlations between the benchmark evaluation accuracy.
Appendix E Evaluation benchmark accuracy and ranking correlation
Figure 11(a) shows the evaluation accuracy scatterplot comparing the reformatted human questions against the synthetic generated benchmark questions for the 5 bolded ensemble models in Table 2. This ensemble produces a Pearson accuracy correlation of $0.745$ between the evaluation accuracies for synthetic and human benchmarks. Each plot marker is the average evaluation accuracy for a given model-dataset combination. While this model ensemble produces reasonable overall accuracy correlation, occasionally specific LMs generate undesirable benchmarks. Figure 12 showcases a failure where the synthetic benchmark generated by gpt-4.1-mini contains overly easy questions where all models perform near $100\%$ evaluation accuracy (Pearson accuracy correlation $0.402$.). This over-performance is likely linked to the question length problem.
Figure 12: Synthetic over-performance for generating model gpt-4.1-mini indicating overly easy questions (Pearson accuracy correlation $0.402$.)
Figure 13 plots the evaluation model performance across all datasets, comparing the synthetic and human accuracy Spearman ranking correlation of the evaluation model capability. This matches the use case of a user selecting between various models for a specific application relevant to the provided grounding documents. Figure 13(a) is the rank order barchart equivalent of Figure 11(a) and Figure 13(b) is the rank order barchart equivalent of the failure case in Figure 12.
(a) Benchmark evaluation accuracy generated by Llama-4-Maverick-17B-128E-Instruct producing a Spearman rank correlation $0.968$.
(b) Benchmark evaluation accuracy generated by gpt-4.1-mini demonstrating over-estimation of LM capability with a Spearman rank correlation $0.564$.
Figure 13: Comparison of benchmark rank ordering between synthetic and human evaluation accuracy.
@@ Appendix E Evaluation benchmark accuracy and ranking correlation
Table 4 showcases the Spearman rank correlation between evaluation benchmarks written by humans and the synthetic data versions. The table includes data for all three of the human question reformatting models used to demonstrate the minimal impact that reformatting has on the overall high level results. Overall, the best outcome is to ensemble together a few questions generated by various reasonably capable LMs to avoid the known slight self-bias/self-preference issue. While we used Llama-4-Maverick as the text feature extractor for measures like groundedness and clarity, that model was unexpectedly dominant in these final results.
Table 4: Spearman rank correlation between synthetic and human benchmarks for multiple choice (MC) and open-ended (OE) evaluations across generating models. Ensemble = bolded models.
@@
Generating Model MC Rank OE Rank MC Rank OE Rank MC Rank OE Rank
Ensemble 0.9679 0.9209 0.9429 0.9253 0.9821 0.8813
Llama-3.2-3B-Instruct 0.6143 0.7143 0.6929 0.7143 0.6536 0.6791
Llama-3.3-70B-Instruct 0.9571 0.8813 0.9643 0.8681 0.9714 0.8242
Llama-3.1-8B-Instruct 0.7964 0.7275 0.7321 0.8110 0.7714 0.7275
Llama-4-Maverick 0.9679 0.9648 0.9536 0.9516 0.9821 0.9209
gemma-3-12b-it 0.8929 0.8198 0.8964 0.8418 0.9000 0.7890
gemma-3-27b-it 0.9571 0.9516 0.9857 0.9473 0.9714 0.9121
gemma-3-4b-it 0.6393 0.9209 0.6571 0.8989 0.6857 0.8637
gpt-4.1 0.9179 0.8549 0.9250 0.8725 0.9357 0.8110
gpt-4.1-mini 0.5643 0.8418 0.6286 0.8549 0.6071 0.7934
o4-mini 0.9250 0.7934 0.9036 0.8374 0.9429 0.7714
phi-4 0.6964 0.1780 0.6250 0.3011 0.7607 0.1736
Qwen2.5-7B-Instruct 0.7321 0.8901 0.7143 0.8945 0.7786 0.8505
Table 5 contains the benchmark evaluation Pearson accuracy correlation between the reformatted human questions and the synthetic data generated questions. Similarly to the Spearman ranking, Llama-4-Maverick performed very well in these tests. However, unlike the ranking data, using a ensemble of models produces significantly worse Pearson accuracy correlation numbers under some circumstances. Likely due to the long and overly easy questions written by gpt-4.1 and o4-mini.
Table 5: Pearson accuracy correlation between synthetic and human benchmarks for multiple choice (MC) and open-ended (OE) evaluations across generating models. Ensemble = bolded models.
@@
Generating Model MC Corr OE Corr MC Corr OE Corr MC Corr OE Corr
Ensemble 0.7452 0.7095 0.7248 0.5630 0.6569 0.4798
@@
Appendix F Question clarity and groundedness
We observed two failure modes with respect to grounding: Either the model hallucinates content absent from and not relevant to the context chunk (phi-4 commonly did this), or the generating model uses knowledge in its weights to expand upon the information presented in the context (o4-mini). For example, phi-4 generated the following question from the provided context:
2
3"context": "Victoria later described her childhood as \"rather melancholy\". Her mother was extremely protective, and Victoria was raised largely isolated from other children under the so-called \"Kensington System\", an elaborate set of rules and protocols devised by the Duchess and her ambitious and domineering comptroller, Sir John Conroy, who was rumoured to be the Duchesss lover. The system prevented the princess from meeting people whom her mother and Conroy deemed undesirable (including most of her fathers family), and was designed to render her weak and dependent upon them. The Duchess avoided the court because she was scandalised by the presence of King Williams bastard children, and perhaps prompted the emergence of Victorian morality by insisting that her daughter avoid any appearance of sexual impropriety. Victoria shared a bedroom with her mother every night, studied with private tutors to a regular timetable, and spent her play-hours with her dolls and her King Charles spaniel, Dash. Her lessons included French, German, Italian, and Latin, but she spoke only English at home."
Appendix G Case study on two academic papers
Our approach was demonstrated on two academic papers: “Current Solutions and Future Trends for Robotic Prosthetic Hands” (labeled annurev-control-071020-104336[44] and “Recent Advances in Large Language Model Benchmarks against Data Contamination: From Static to Dynamic Evaluation” (labeled arXiv_2502_17521v1[45]. Figure 14 showcases LM benchmark evaluation performance on these two datasets; the MCQ results show most models performing well before a sharp dropoff with 4B and fewer parameter models and the Open-Ended results with a much smoother decline in performance. Multiple-choice questions can be easier to answer due to the presence of the answer options, even if the evaluation framework shuffles the option order. This effect is likely the cause of the MCQ evaluation accuracy to remaining higher for longer.
@@ Appendix G Case study on two academic papers
Figure 14: Evaluation benchmark rank ordering comparing MCQ to Open-Ended questions for two papers drawn from current research work [44, 45].
The following example questions were drawn from the MCQ synthetic benchmark constructed from arXiv_2502_17521v1 [45].
@@
7 "D": "Temporal cutoff generates data points using LLMs, while rule-based generation uses manual curation."},
8
9"answer": "A"},
10
11{"question": "What is a potential future research direction for improving the design and standardization of dynamic benchmarking methods for Large Language Models (LLMs), given the current limitations and proposed evaluation criteria?",
12
13"choices": {
@@
17 "D": "Focusing solely on post-hoc contamination detection methods."},
18
19"answer": "A"}
@@
2
3"answer": "By leading to misleading conclusions that favor contaminated models."},
4
5{"question": "What are the implications of using LLMs to rewrite benchmark samples while preserving their original difficulty levels, as seen in methodologies like ITD, on the overall diversity and complexity of the resulting benchmark datasets?",
6
7"answer": "The resulting datasets may maintain complexity but risk lacking diversity if not combined with other rewriting strategies."}